# Large Language Model

### v1

@@ 5.3 Qualitative results
Table 4: Exploration experiments of using the untraining EventVL initialized from InternVL2-2B [7]. “wo ev./im.” denotes that we don’t use the event or image modality as input.
@@
+N-ImageNet N-Caltech101 - 0.435 0.289 0.636 1.273
w/o im. 0.180 0.162 0.444 0.377
[…]
w/o im. 0.031 0.126 0.243 0.089
[…]
w/o im. 0.121 0.151 0.386 0.889
+N-ImageNet +HARDVS N-Caltech101 - 0.425 0.287 0.634 1.275
w/o im. 0.176 0.161 0.445 0.372
[…]
w/o im. 0.386 0.317 0.618 2.522
[…]
w/o im. 0.117 0.148 0.379 0.833
+N-ImageNet +HARDVS +DSEC N-Caltech101 - 0.460 0.295 0.638 1.480
w/o im. 0.205 0.167 0.462 0.573
[…]
w/o im. 0.396 0.314 0.628 2.603
[…]
w/o im. 0.115 0.145 0.377 0.823
Table 5: Ablation study about the processed dataset choice for training by using event data as input. Note that categories from training sets aren’t overlapped with the test sets. “w/o im.” denotes that we don’t use the image modality as input.
5.4 Ablation Study
In this section, we utilize the zero-shot caption generation evaluation setting to perform the comprehensive ablation studies for investigating each effective module. More ablation studies can be found in the appendix.
Dataset Quality Assessment. In this part, we investigate the effectiveness of our proposed dataset. As shown in Table 4, we directly utilize the untraining EventVL initialized from InternVL2-2B to test on all validation sets for observing the current SOTA MLLM weights on the processed dataset performance. Experiments show that event representations transferred in the red-blue colour map make the MLLM confused, leading to serious understanding degradation. Furthermore, event data hinders the MLLM’s understanding of images. These observations prove event-based understanding deficiency of MLLMs and the urgency of our motivation. Meanwhile, we explore the enhancement of model training performance as shown in Table 5. We can get the conclusion that the single domain dataset training only makes a positive effect on the corresponding val sets while getting a tiny improvement on other val sets.
5.5 Extensive Application
@@
Table 6: Event retrieval results on processed N-Caltech101 [31].
We also explore other downstream applications based on our framework. For example, EventVL can be easily re-developed for event-image-text retrieval tasks. We utilize EventBind [52] for comparison as shown in Table 6. Experiment settings and discussion can be found in the appendix.
6 Conclusion
In this work, we introduce EventVL, the first event-based MLLM designed to achieve comprehensive understanding of event stream data. Through the integration of an advanced event encoder that captures rich semantic information, the novel Event Spatiotemporal Representation for efficient feature aggregation, and the Dynamic Semantic Alignment mechanism for fine-grained feature alignment, EventVL significantly advances the understanding and reasoning capabilities for event-based data. Additionally, we contribute a large-scale, high-quality annotated dataset with nearly 1.4 million event-image-text pairs, providing a robust foundation for future event-based research. We believe that EventVL sets a new direction for chat-centric event stream comprehension and lays the groundwork for future breakthroughs in event-driven multimodal systems.
Supplementary Material
7 Related Work
Event-based Vision. The microsecond-level temporal resolution, high dynamic range (typically 120 dB compared to the 60 dB of standard cameras), and power efficiency of event cameras represent a paradigm shift in imaging, moving beyond traditional frame-based methods [14]. Numerous event-based low-level imaging tasks, including recognition, perception, localization, and reconstruction, have been developed, covering areas such as object recognition [11, 33], object detection [44, 15], optical flow [12, 27], semantic segmentation [25], depth estimation [32], and object reconstruction [49, 22, 42, 13, 21], often utilizing techniques like NERF or Gaussian Splatting. While event-based models have demonstrated success in these traditional perception tasks, the exploration of event-based MLLMs (Multimodal Large Language Models) remains largely uncharted. In this paper, we focus on a novel application—event captioning and description for comprehensive event stream understanding. This endeavor aims to address the challenges posed by sparse, asynchronous, and high-temporal-resolution event data, with the goal of generating explicit language descriptions. Such capabilities are particularly crucial for safety-critical applications, such as drone or in-vehicle perception systems.
Figure 5: Problem list for different kinds of datasets. “Static Images” is for N-ImageNet [23] and N-Caltech101 [31], “Human Motions” is for [40] and “Drive Scenes” is for DSEC [16].
8 Preliminaries
The original stream output of an event camera consists of a set of asynchronously generated events, where each event is represented by $(x,y,t,p)$. Here, $(x,y)$ are the spatial coordinates, $t$ is the timestamp, and $p$ indicates the polarity of the intensity change: $+1$ for increased brightness and $-1$ for decreased brightness. These events feature a high dynamic range greater than 120 dB and a high temporal resolution equivalent to thousands of frames per second. However, directly processing event streams is challenging [14]. Previous works often represent events in frame-like representation, such as gray-style [43] or red-blue color map [25, 52], to simplify cross-modal alignment.
9 Data Engineering Setting
Algorithm 1 Generate Adaptive Ratios
Input: $n_{\text{min}},n_{\text{max}}$ are the min./max. number of patch.
Output: $\mathcal{K}$ is the adaptive ratio set.
1:$\mathcal{K}\leftarrow\{\}$
2:for $n\leftarrow n_{\text{min}}$ to $n_{\text{max}}$ do
3:    for $i\leftarrow 1$ to $n$ do
4:       for $j\leftarrow 1$ to $n$ do
5:          if $i\times j\leq n_{\text{max}}$ and $i\times j\geq n_{\text{min}}$ do
6:             $\mathcal{K}\leftarrow\mathcal{K}\cup\{(i,j)\}$
[…]
8:return $\mathcal{K}$
In this section, we mainly introduce the origin dataset information, the problem list and process details.
9.1 Origin Dataset
N-ImageNet [23]. As the largest event camera dataset, it contains 1.78 million event streams and 1,000 classes. The data were captured with a moving 480 × 640 resolution Samsung DVS Gen3 event camera under the same environmental conditions. N-ImageNet also provides variants of the test set captured with different camera motions and brightness, serving as a benchmark to evaluate the robustness of event classifiers. The authors create 9 variants of the same test set. Variants 1-5 change the camera motions used to trigger events, where different moving directions, frequencies, and amplitudes of the camera trajectory are employed. Variants 6-9 alternate the lighting conditions of the environment, such as extremely low or high illuminations. Overall, these variations cause a large degradation in the performance of existing event-based classifiers.
N-Caltech101 [31]. N-Caltech101 contains 8,246 samples from 101 classes, recorded by a moving 180 × 240 resolution ATIS system in front of a monitor displaying still images from the original RGB Caltech101 dataset.
HARDVS [40]. HARDVS is the first large-scale realistic neuromorphic human activity recognition dataset. Specifically, the HARDVS dataset contains more than 100K video clips recorded with a DAVIS346 camera, each of them lasting for about 5-10 seconds. It contains 300 categories of human activities in daily life, such as drinking, riding a bike, sitting down, and washing hands. The following factors are taken into account to make HARDVS data more diverse, including multi-views, illuminations, motion speed, dynamic background, occlusion, flashing light, and photographic distance.
DSEC  [^1] https://dsec.ifi.uzh.ch/ [16]. DSEC is an extensive dataset designed for advanced driver-assistance systems (ADAS) and autonomous driving research, with a particular focus on event-based vision and stereo vision. Different from DDD17 [3], DSEC dataset combines data from event-based cameras and traditional RGB cameras. Including event-based cameras (which capture changes in light intensity) alongside regular cameras provides a rich, complementary data source for perception tasks. The dataset typically features high-resolution images and event data, providing detailed visual information from various driving conditions, including urban, suburban, and highway environments, various weather conditions, and different times of the day. This diversity is crucial for developing systems that can operate reliably in real-world conditions. In this paper, we only utilize the training set of the DSEC for post-processing. In detail, we first split each scene video into multiple segments by 24 frames per segments. Then we utilize these segment for generating our description by Section 3.
9.2 Processing details
Our different problem lists are as shown in Figure 5. In each round annotating the data, we randomly sample the problem from the corresponding lists. For N-ImageNet, we directly utilize the whole event-pairs dataset for training, and we do not use its variants. For N-Caltech101, we utilize it only to validate the effectiveness of our proposed EventVL. For HARDVS, we directly utilize its red-blue color map representation instead of the raw event data for efficiency. In each category, we uniformly sample 14 frames for each video for post-processing. If videos contain less than 14 frames, we directly utilize the whole video. For DSEC, we only sample 14 frames in each segment for training efficiency.
Figure 6: Case visualization of the proposed Event Spatiotemporal Representation.
10 Methods
As shown in Figure 6, we give a case to visualize the event stream processed by our proposed Event Spatiotemporal Representation. We decouple the event stream from time and space to obtain a comprehensive representation of the event stream. Then, We give more explanation of the proposed Adaptive Spatial Split. For predefined set $\mathcal{K}$, we utilize Algorithm 1 to complete the $\mathcal{K}$. As we set $n_{min}$ to 1 and set $n_{max}$ to 6, we have $\mathcal{K}\in${1:1,1:2,…,6:1}. Then, we utilize the $\mathcal{K}$ for following event spatial segmentation in Section 4.1.
11 Experiments
11.1 Implementation
Fundamental Module. We selected the InternViT-300M-448px  [^2] https://huggingface.co/OpenGVLab/InternViT-300M-448px [7] as the foundational architecture for both the image and event encoders, given its proven effectiveness in various image and video understanding benchmarks. This model contains 24 attention blocks, as shown in Figure 7. The term “Drop” refers to the drop path, commonly applied in the main path of residual blocks. In our setup, $B$ denotes the batch size, $C$ represents the channel number, and $H$ and $W$ are the height and width of the image or event. $P$ is the patch number, and $C_{\phi}$ is the final output embedding channel number. We set $H=W=448$, $C_{\phi}=2048$, and $P=256$. We also leverage the pretrained weights of InternViT for initialization, allowing us to transfer spatial prior knowledge from images to events, which accelerates training and improves data efficiency for EventVL. For the text decoder, we chose InternLM2-1.8B [4], which has shown significant improvements in capabilities such as reasoning and coding.
Figure 8: Qualitative results about processed zero-shot caption generation by utilizing N-ImageNet [23] val set. Red color denotes model wrong description and orange color denotes the accurate object description. Underline text means redundant description.
General Setting. In Dynamic Semantic Alignment, we set $n_{min}=1$ and $n_{max}=6$. We also set the $\lambda_{1}=1$ and $\lambda_{2}=1$ for balancing the Multimodal alignment. Caused by the difference between the event number scale between N-ImageNet and N-Caltech101, we set different event point numbers per frame $n_{\epsilon}$ for each dataset, which we set $n_{\epsilon}=40000$ for N-ImageNet and $n_{\epsilon}=20000$ for N-Caltech101. If the total event point number $N_{\epsilon}$ does not fulfill the demand of all level event-based representation from Section 4.1, we pad this event stream for post-processing. Note that we do not utilize Event Spatiotemporal Representation for video-event pairs from HARDVS [40] and DSEC [16] datasets, which contain enough temporal and spatial information. We directly utilize cosine similarity from Equation 1 to align the video-event modalities. The final optimization objectives still are the same as the image-event pairs datasets training. Due to the unbalanced data scales between these domains, we adopt the weighted sampling strategy for hybrid Training. We set the training sampling weight of 0.6 for N-ImageNet, 0.1 for DSEC, and 0.3 for HARDVS. COCO caption metrics [6] is used for quantitative analysis model performance. All training is performed on eight Nvidia GTX A6000 GPUs for approximately 120 hours. We set the batch size to 2 and adopt the DeepSpeed  [^3] https://github.com/microsoft/DeepSpeed ZeRO-1 training strategy for training efficiency.
Zero-shot Setting. We utilized N-ImageNet, HARDVS and DSEC train splits for training. To overcome overfitting, we adopt 1 epoch for training. We evaluate EventVL on the val sets for zero-shot description generation with only event data as input.
Few-shot Setting. For few-shot experiments, we sample $n$-shot data from the all data of N-ImageNet, HARDVS, and DSEC for few-shot training. Then, we utilize the remaining datasets including N-Caltech101 to evaluate its performance. We adopt 3 epochs to perform all few-shot training in order to learn enough event-based category knowledge.
L1 L2 L3 N-Caltech101 HARDVS N-ImageNet
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
- - 0.325 0.769 0.582 1.885 0.434 0.548
- 0.349 0.789 0.598 2.409 0.445 0.554
- 0.353 0.786 0.601 2.411 0.441 0.557
0.377 0.823 0.628 2.603 0.462 0.573
Table 7: Different level of the event-data aggregation choice for event batch construction. “L1/L2/L3” denotes the Level-1/2/3.
11.2 More Cases For Zero-shot Evaluation
Figure 9: Qualitative results about processed zero-shot caption generation by utilizing HARDVS [40] val set. Red color denotes model wrong description and orange color denotes the accurate object description.
As shown in Figure 8 and Figure 9, we provide additional examples to validate the superiority of our proposed EventVL framework using only event data as input. These examples demonstrate that EventVL captures more fine-grained details from the given event stream compared to VideoLLaMA2, highlighting the effectiveness of our framework.
Figure 10: Qualitative results about multi-round dialogue by utilizing N-ImageNet [23] val set. Blue color denotes the detailed description.
11.3 Multi-round Dialogue
As illustrated in Figure 10, we present several examples demonstrating the application of our proposed EventVL framework in multi-round dialogue based on event data. These examples reveal that EventVL can understand comprehensive semantics within event data while also capturing finer object details, such as describing “a clear blue sky”.
$n_{max}$ N-Caltech101 HARDVS N-ImageNet
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
1 0.310 0.756 0.565 2.398 0.431 0.528
5 0.341 0.810 0.603 2.575 0.459 0.569
6 0.377 0.823 0.628 2.603 0.462 0.573
7 0.388 0.841 0.633 2.648 0.473 0.592
8 0.386 0.843 0.635 2.649 0.468 0.575
Table 8: Different $n_{max}$ settings for Dynamic Semantic Alignment. Note that we default the $n_{min}=1$.
11.4 Ablation Study
Dataset Quality Assessment. We investigate the impact of training on individual datasets to enhance model performance, as summarized in Table 5. Our findings indicate that training across multiple domains improves the model’s general event-based understanding by mitigating overfitting issues seen in single-domain training. For example, this approach yields nearly a 20% increase in CIDEr on the N-Caltech101 dataset. Additionally, incorporating scene data (DSEC) enhances the model’s ability to capture broader contextual details, evidenced by CIDEr gains on HARDVS compared to training without DSEC. Moreover, improvements in event-based understanding bolster the model’s overall multimodal comprehension. Compared to results in Table 4, our method achieves 1.769 CIDEr (+46.3%) on the N-ImageNet validation set, significantly outperforming the untrained model (1.306 CIDEr).
@@ 11.4 Ablation Study
Event Representation. As shown in Table 8, we observe that excluding the dynamic patch setting ($n_{max}=1$) results in significant performance degradation. Lower values of $n_{max}$, such as $n_{max}=5$, lead to coarse alignment between images and events due to insufficient semantic information. Conversely, increasing $n_{max}$ to values like $n_{max}=8$ does not enhance EventVL’s understanding, as its category knowledge reaches inherent limits. Additionally, larger values of $n_{max}$ consume more CUDA memory. To balance performance and resource consumption, we select $n_{max}=6$ for efficient temporal interaction and semantic extraction. We also examine various levels of aggregation in the Adaptive Spatial Split process. As shown in Table 7, results indicate that all-level event aggregation provides the best performance, while any omission of representation at any level causes a decline in our framework’s effectiveness.
Optimization Setting. As shown in Table 9, we first analyze the main components of our training objectives. When using only cross-entropy for event-text alignment, the model performance decreases by 9.7% in CIDEr on the N-ImageNet evaluation. Incorporating the image prior improves performance, likely by bridging semantic gaps between text and event data. The proposed $\mathcal{L}_{c}$ introduces an additional 5.8% improvement in CIDEr on N-ImageNet, enhancing multimodal alignment.
We also examine the impact of varying weight settings in the optimization objectives of Equation 5. As shown in Table 10, the results indicate that text-based alignment exerts a greater influence on the model’s understanding of event data. Lower values of $\lambda_{1}$ correlate with a more significant drop in performance, demonstrating its importance in our optimization strategy.
Prior $\mathcal{L}_{c}$ N-Caltech101 HARDVS N-ImageNet
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
- - 0.315 0.725 0.524 1.996 0.324 0.466
- 0.342 0.747 0.568 2.394 0.358 0.491
- 0.356 0.734 0.545 2.373 0.334 0.524
0.377 0.823 0.628 2.603 0.462 0.573
Table 9: Various loss construction for training objective Equation 5. “Prior” denotes the image prior and $\mathcal{L}_{c}$ denotes the image-event embedding similarity loss.
11.5 Failure Cases
$\lambda_{1}$ $\lambda_{2}$ N-Caltech101 HARDVS N-ImageNet
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
0.2 1 0.298 0.368 0.512 1.235 0.324 0.398
0.5 1 0.319 0.440 0.553 1.549 0.330 0.435
1 0.1 0.291 0.667 0.467 1.874 0.379 0.317
1 0.2 0.298 0.689 0.488 2.102 0.387 0.341
1 0.5 0.345 0.724 0.574 2.263 0.413 0.451
1 1 0.377 0.823 0.628 2.603 0.462 0.573
Table 10: Different $\lambda_{1}$ and $\lambda_{2}$ setting for the training objective of Equation 5.
As shown in Figure 11, we present several failure cases of our proposed EventVL. We summarize three types of errors that occur in EventVL as follows: Misclassification (Case (a,f,g)), Misinterpretation (Case (d,e,i)) and Recognition Failures (Case (b,c,h)). For Misclassification, our model limits in distinguishing between different animal categories when facing the poor event stream quality. For instance, in case (a), the model misidentifies a chicken as “a dog”. Second, for Misinterpretation, our model fails to accurately recognize ongoing activities or scene elements such as case (d), where an elderly person playing a harp is mistakenly described as “painting a window”. Even though multimodal alignment forces the model to learn such knowledge in post-processing, the model still falls short in capturing event data surface details such as object textures. Finally, we also observe the Recognition Failures. For example, in case (c), the model fails to mention the prominent pink flowers in the scene. These failure cases reveal a fundamental challenge in applying vision-based models to event camera data. The distinctive texture patterns and unique visual representations generated by event cameras can introduce ambiguity that conventional vision models struggle to interpret accurately. In the future, we will explore more efficient strategy to bridge this gap between images and events for better event-based understanding.
11.6 Event Retrieval
The complex architecture of decoder-only LLMs makes it challenging to adapt them as text encoders. Recent research has attempted to modify the inherent mechanisms of LLMs to handle text representation tasks [2], but this approach often lacks efficiency. Inspired by [20], we instead freeze the overall model and add trainable attention layers to both the encoder and decoder, followed by contrastive learning. This approach allows for efficient retrieval tasks with strong performance. Specifically, we first add an attention layer to each frozen encoder, and a linear projection layer to the text decoder. These additional layers are trainable, designed to reduce the high-dimensional complexity of the model to a lower-dimensional space for more efficient alignment. We then apply contrastive learning [5] to achieve significant semantic alignment across modalities. We set the training epoch to 3, keeping other parameters consistent with previous experiments. As shown in Table 6, our proposed strategy achieves comparable performance to EventBind, demonstrating the extensibility of our EventVL framework.
12 Discussion

### v2

@@ D-C Ablation Study
TABLE XII: Ablation study about the processed dataset choice for training by using event data as input. Note that categories from training sets aren’t overlapped with the test sets. “w/o im.” denotes that we don’t use the image modality as input.
@@
+N-ImageNet N-Caltech101 - 0.445 0.295 0.649 1.302
w/o im. 0.184 0.165 0.453 0.385
[1pt/1pt] HARDVS - 0.187 0.239 0.482 1.010
w/o im. 0.032 0.129 0.248 0.091
[1pt/1pt] N-ImageNet - 0.212 0.224 0.495 1.789
w/o im. 0.123 0.154 0.394 0.907
+N-ImageNet
+HARDVS
N-Caltech101 - 0.434 0.293 0.647 1.301
w/o im. 0.179 0.164 0.454 0.380
[1pt/1pt] HARDVS - 0.445 0.349 0.669 3.015
w/o im. 0.394 0.323 0.630 2.572
[1pt/1pt] N-ImageNet - 0.210 0.224 0.495 1.791
w/o im. 0.119 0.151 0.387 0.850
+N-ImageNet
+HARDVS
+DSEC
N-Caltech101 - 0.469 0.301 0.651 1.510
w/o im. 0.223 0.183 0.477 0.599
[1pt/1pt] HARDVS - 0.445 0.347 0.670 3.055
w/o im. 0.407 0.325 0.639 2.619
[1pt/1pt] N-ImageNet - 0.210 0.224 0.495 1.805
w/o im. 0.125 0.157 0.397 0.853
D-D Multi-round Dialogue
As illustrated in Figure 4 and 11, we present several examples showcasing the application of our proposed EventVL framework in multi-round dialogue using event data. These examples demonstrate that EventVL not only understands the comprehensive semantics within event data but also captures finer object details, like describing “a clear blue sky”.
Figure 11: Qualitative results about multi-round dialogue by utilizing HARDVS [16] val set.
D-E Extensive Application
We explore diverse downstream applications via EventVL.
TABLE XIII: Event retrieval results on processed N-Caltech101 [26].
@@ D-E Extensive Application
Retrieval Tasks. EventVL can be readily adapted for event-image-text retrieval tasks. For comparison, we employ EventBind [7], as shown in Table XIII. Drawing inspiration from [49], we adopt a parameter-efficient approach by freezing the core model and introducing trainable attention layers in both the encoder and decoder, followed by contrastive learning. This strategy enables efficient retrieval while maintaining strong performance. Specifically, we integrate an additional attention layer into each frozen encoder and introduce a linear projection layer in the text decoder. These trainable components serve to reduce the high-dimensional complexity of the model, facilitating a more efficient alignment between modalities. We then apply contrastive learning [50] to enhance semantic alignment across different modalities. The model is trained for three epochs, with all other hyperparameters kept consistent with previous experiments. As reported in Table XIII, our proposed adaptation achieves performance comparable to EventBind, highlighting the extensibility of the EventVL framework. However, our approach exhibits slightly lower performance, likely due to the inherent challenges of adapting decoder-only large language models (LLMs) as text encoders [51]. The complex architecture of decoder-based LLMs can lead to suboptimal text embeddings, ultimately affecting retrieval accuracy.
Figure 12: Illustration of the detection by integrating EventVL and GroundingDINO [6]. Our EventVL generate the rich description for specified object road and trees for open-set object detection.
Figure 13: Illustration of the detection by integrating EventVL and GroundingSAM [52]. Our EventVL generate the rich description for specified object lamppost for open-set segmentation.
Detection. To assess the generalization capability of EventVL in object-level understanding and reasoning, we integrate it with GroundingDINO, leveraging its open-set object detection capabilities. Given an input query, EventVL first generates textual descriptions of potential objects present in the event streams. These descriptions are then utilized by GroundingDINO [6] to localize the corresponding objects within the data. As illustrated in Figure 12, EventVL exhibits strong object-level reasoning abilities, accurately identifying and localizing target objects such as tree and road within the scene. This synergy between EventVL and GroundingDINO highlights its robust inferential object detection capabilities and its potential for applications in domains such as autonomous driving, particularly in challenging environments like low-light.
Segmentation. To validate the scalability of EventVL in complex dense prediction tasks, we integrate it with GroundingSAM [52] for event-stream segmentation. EventVL provides semantic cues that guide GroundingSAM in generating precise instance masks, enabling accurate segmentation of individual objects in the scene, such as the lamppost. The qualitative results in Figure 13 demonstrate the effectiveness of our approach in achieving high-precision segmentation under the challenging conditions inherent to event-based data. This highlights EventVL adaptability and potential for fine-grained scene understanding.
D-F Failure Cases
As shown in Figure 14, we present several failure cases of our proposed EventVL. We summarize three types of errors that occur in EventVL as follows: Misclassification (Case (a,f,g)), Misinterpretation (Case (d,e)) and Recognition Failures (Case (b,c,h)). For Misclassification, our model limits in distinguishing between different animal categories when facing the poor event stream quality. For instance, in case (a), the model misidentifies a chicken as “a dog”. Second, for Misinterpretation, our model fails to accurately recognize ongoing activities or scene elements such as case (d), where an elderly person playing a harp is mistakenly described as “painting a window”. Even though multimodal alignment forces the model to learn such knowledge in post-processing, the model still falls short in capturing event data surface details such as object textures. Finally, we also observe the Recognition Failures. For example, in case (c), the model fails to mention the prominent pink flowers in the scene. These failure cases reveal a fundamental challenge in applying vision-based models to event camera data. The distinctive texture patterns and unique visual representations generated by event cameras can introduce ambiguity that conventional vision models struggle to interpret accurately. In the future, we will explore more efficient strategy to bridge this gap between images and events for better event-based understanding.
@@ D-F Failure Cases
Appendix E Discussion